# A kernel conditional two-sample test

### v1

@@ Appendix B Bootstrapping
Case 1 in Theorem 4.3 generalizes Corollary 3.4 in Abbasi-Yadkori [12] to separable-Hilbert-space-valued outputs without introducing conservatism, as it recovers exactly the same bound as in this reference under the assumptions there. It also generalizes Theorem 1 in Chowdhury and Gopalan [20] by removing the regularity assumptions connecting the input set $\mathcal{X}$ and the kernel $K$ (in the reference, $\mathcal{X}$ is compact and $K$ is bounded). It further generalizes this same result in yet another way by removing the assumption that $K$ should be trace-class, replacing it by the (significantly milder) assumption that ${K_{0}}$ should be so and hereby allowing $K=k\cdot\mathrm{id}_{\mathcal{G}}$ as a special case. To the best of our knowledge, this is the first time-uniform uncertainty bound for KRR allowing this essential case.
@@
The assumption that the kernel $K$ is UBD in Case 1 above is the central methodological contribution of the proof. It enables us to generalize the proof method of Abbasi-Yadkori [12] by leveraging a tensor factorization of the RKHS $\mathcal{H}_{K}$ (Proposition D.12). Specifically, Proposition D.12 allows avoiding the usual problems encountered with non-trace-class kernels — more precisely, that trace-dependent quantities explode — by replacing them with the corresponding quantities for the elementary block ${K_{0}}$. Intuitively, all of the “estimation machinery” is performed in the RKHS of ${K_{0}}$, which is a trace-class kernel and for which classical methods work well, and we simply need to ensure that the noise is “compatible” with this structure. Here, this compatibility takes the form of the operator $R_{\mathcal{V}}$ being trace-class. Interestingly, a special case of Proposition D.12 has been used in Li et al. [40] (Theorem 1) to identify $\mathcal{H}_{K}$ with a space of Hilbert-Schmidt operators in the case $K=k\cdot\mathrm{id}_{\mathcal{G}}$, hereby avoiding problems with the trace of $K$ by working with that of $k$, similarly to us. It appears that many arguments of Li et al. [40] also hold under the more general assumption that $K$ is UBD. An interesting avenue for future work is thus to investigate to what extent this generalization carries over, and whether the derived learning rates [40, Theorem 2] hold under UBD. Summarizing, we believe the UBD assumption and the resulting structure of the RKHS identified in Proposition D.12 is a relevant new idea to handle a large class of non-trace-class kernels and can be used in other areas of kernel methods such as learning theory with loss functions compatible with the structure.
@@
Appendix D Generalized confidence bounds in vector-valued least squares
We provide in this section the proof of Theorem 4.3. Specifically, we show a more general result which does not assume that the space $\mathcal{H}_{K}$ in which the estimator lies is \@iaciRKHS RKHS; only that it is a generic Hilbert space $\mathcal{H}$. The result we show is given later in this section in Theorem D.23 and pertains to confidence bounds in vector-valued least squares estimation under general assumptions on the sampling operators; KRR is then obtained for a specific choice of the measurement operators and Hilbert space where estimation is performed.
This section is organized as follows. We first provide in Section D.1 the general setup of vector-valued least squares estimation, and identify that the main challenge for confidence bounds is bounding a specific term involving the noise. We then provide in Sections D.2 and D.3 methods to bound this term under different assumptions that we formalize. The main intermediate results of those sections are Lemma D.4, Theorem D.14, and Corollary D.20, and are essential in the proof of Theorem D.23, which we only state and prove in Section D.4 as we provide a gentle introduction to its assumptions first. We finally leverage it in Section D.5 to show Theorem 4.3.
D.1 Setup and goal
@@
$\hskip 28.45274pt+\left\|\left(\sum_{n=1}^{N}L_{n}^{\ast}L_{n}+\lambda\mathrm{id}_{\mathcal{H}}\right)^{-1}\sum_{n=1}^{N}L_{n}^{\ast}y_{n}\right\|_{\sum_{n=1}^{N}L_{n}^{\ast}L_{n}+\mathrm{id}_{\mathcal{H}}}^{2}-\left\|\left(\sum_{n=1}^{N}L_{n}^{\ast}L_{n}+\lambda\mathrm{id}_{\mathcal{H}}\right)^{-1}\sum_{n=1}^{N}L_{n}^{\ast}y_{n}\right\|_{\sum_{n=1}^{N}L_{n}^{\ast}L_{n}+\lambda\mathrm{id}_{\mathcal{H}}}^{2}$
@@
$\hskip 14.22636pt=\left\|h-\left(\sum_{n=1}^{N}L_{n}^{\ast}L_{n}+\lambda\mathrm{id}_{\mathcal{H}}\right)^{-1}\sum_{n=1}^{N}L_{n}^{\ast}y_{n}\right\|_{\sum_{n=1}^{N}L_{n}^{\ast}L_{n}+\lambda\mathrm{id}_{\mathcal{H}}}^{2}+\text{Terms without $h$}$
@@
Our general goal is to obtain a finite-sample confidence region for the estimator $h_{N,\lambda}$ of $h^{\star}$. While we only state the precise result in Section D.4, we want — in short — a high-probability bound on the quantity $\lVert\bar{L}h_{N,\lambda}-\bar{L}h^{\star}\rVert_{\bar{\mathcal{G}}}$. Here $\bar{\mathcal{G}}$ is a Hilbert space and $\bar{L}\in\mathcal{L}_{\mathrm{b}}(\mathcal{H},\bar{\mathcal{G}})$ is a bounded operator capturing an appropriate property we are interested in; they are degrees of freedom of the problem. We begin with the following observation. This result, which is well-known in the case of KRR (see for example Abbasi-Yadkori [12, Chapter 3] or Chowdhury and Gopalan [19]), results from elementary algebraic manipulations, the triangle inequality, and properties of the operator norm.
Lemma D.2.
@@
$\begin{split}\lVert\bar{L}(M_{N}^{\star}M_{N}+\lambda\mathrm{id}_{\mathcal{H}})^{-1/2}\rVert_{\mathcal{L}_{\mathrm{b}}(\mathcal{H},\bar{\mathcal{G}})}&=\sqrt{\left\lVert\bar{L}(M_{N}^{\star}M_{N}+\lambda\mathrm{id}_{\mathcal{H}})^{-1/2}\left(\bar{L}(M_{N}^{\star}M_{N}+\lambda\mathrm{id}_{\mathcal{H}})^{-1/2}\right)^{\star}\right\rVert_{\mathcal{L}_{\mathrm{b}}(\mathcal{H})}}\\ &=\sqrt{\frac{1}{\lambda}\left\lVert\bar{L}\left(\mathrm{id}_{\mathcal{H}}-(M_{N}^{\star}M_{N}+\lambda\mathrm{id}_{\mathcal{H}})^{-1}M_{N}^{\star}M_{N}\right)\bar{L}^{\star}\right\rVert_{\mathcal{L}_{\mathrm{b}}(\mathcal{H})}}.\end{split}$ (21)
@@
D.2 Noise bound for deterministic sampling
@@
Assumption D.3 (Deterministic sampling).
@@
D.3 Noise bound for online sampling with coefficients-transforming sampling operators
@@
Assumption D.5 (Online sampling).
@@
We address this correlation by showing a self-normalized concentration inequality for $\lVert S_{N}\rVert_{(M_{N}^{\star}M_{N}+\lambda\mathrm{id}_{\mathcal{H}})^{-1}}$. The inequality is also time-uniform, meaning that the high-probability set on which the bound holds does not depend on the time $N\in{\mathbb{N}_{>0}}$. It is a generalization of Theorem 3.4 (and Corollary 3.6) in Abbasi-Yadkori [12] from $\mathcal{G}=\mathbb{R}$ to a general separable Hilbert space $\mathcal{G}$. These results are based on the method of mixtures [43]. While we mostly follow the proof strategy from Abbasi-Yadkori [12], we introduce a general structural assumption on the sampling operators (Definition D.11 and Proposition D.12) to handle the case where the sampling operators are not trace-class (in the RKHS case, the case where the kernel is not trace-class). Furthermore, in contrast to Abbasi-Yadkori [12] we do not employ an explicit stopping-time construction to obtain time uniformity, but rather use Ville’s inequality as in Whitehouse et al. [21].
D.3.1 Challenge and goal
@@
Our proof is based on the method of mixtures of Abbasi-Yadkori [12]. A close look at the proof of Theorem 3.4 in the reference reveals that the problematic determinant (24) arises when integrating the function $G_{N}$ (in the notations of the reference) against a Gaussian measure on $\mathcal{H}_{K}$. We circumvent this integration by leveraging a factorization of $\mathcal{H}_{K}$ as a tensor space. While such a factorization $\mathcal{H}_{K}\cong H_{1}\otimes H_{2}$ trivially exists for choices of separable $H_{1}$ and $H_{2}$ based on dimensional arguments (i.e., $\dim(\mathcal{H}_{K})=\dim(H_{1})\cdot\dim(H_{2})$), we choose one in which all of the sampling operators $L_{n}$, $n\in{\mathbb{N}_{>0}}$, admit themselves the simple form $L_{n}\cong\tilde{L}_{n}\otimes\mathrm{id}_{H_{2}}$, and for which $\tilde{L}_{n}$ has nice properties (specifically, $\tilde{L}_{n}^{\star}\tilde{L}_{n}$ is trace-class). In the case where $K=k\cdot\mathrm{id}_{\mathcal{G}}$ and $L_{n}=K(\cdot,x_{n})$, this construction is known and natural: the RKHS $\mathcal{H}_{K}$ is the (closure of the) span of vectors of the form $f\cdot g$, where $f\in\mathcal{H}_{k}$ and $g\in\mathcal{G}$; see for instance Paulsen and Raghupathi [36, Exercise 6.3]. Then, for any $h=\sum_{m=1}^{\infty}f_{m}\cdot g_{m}\in\mathcal{H}_{k\cdot\mathrm{id}_{\mathcal{G}}}$, introducing $h_{\otimes}=\sum_{m=1}^{\infty}f_{m}\otimes g_{m}\in\mathcal{H}_{{K_{0}}}\otimes\mathcal{G}$,
@@
In the more general case where $K$ is only UBD instead of diagonal, we prove that such a factorization also exists (Proposition D.12). In any case, this factorization allows us to write $\mathcal{H}_{K}\cong\mathcal{H}_{{K_{0}}}\otimes\mathcal{V}$. This enables evaluating (a modification of) the function $G_{N}$ on pure tensors $\tilde{h}\otimes v$, where $\tilde{h}\in\mathcal{H}_{{K_{0}}}$ and $v\in\mathcal{V}$. We then leverage a double mixture argument: one for the variable $\tilde{h}$, and one for the variable $v$. The one on $\tilde{h}$ is the same as in Abbasi-Yadkori [12]: a Gaussian measure on $\mathcal{H}_{{K_{0}}}$, which outputs the desired determinant involving ${K_{0}}$. The one on $v$ is not Gaussian and enables recovering the norm we wish to bound, $\lVert S_{N}\rVert_{(V_{N}+\lambda\mathrm{id}_{\mathcal{H}_{K}})^{-1}}$ (up to some regularization term), when combined with Jensen’s inequality. Our double mixture strategy is fundamentally different from what is used for Chowdhury and Gopalan [19, Theorem 1], since we use two different types of mixture distributions, and the second mixture enters in a multiplicative manner (instead of an additive one). We can then conclude using standard arguments involving Ville’s inequality for positive supermartingales, similarly to Whitehouse et al. [21]. Summarizing, this method enables showing the following result, which is a specialization of the main outcome of this section to the case $\mathcal{H}=\mathcal{H}_{K}$ and $L_{n}=K(\cdot,x_{n})^{\star}$.
Theorem D.6 (Theorem D.14, RKHS case).
Under the notations of Sections D.1 and D.3.1, assume that $\mathcal{H}=\mathcal{H}_{K}$, where $K$ is UBD with block decomposition ${K_{0}}:\mathcal{X}\times\mathcal{X}\to\tilde{\mathcal{G}}$ and isometry $\iota_{\mathcal{G}}\in\mathcal{L}_{\mathrm{b}}(\tilde{\mathcal{G}}\otimes\mathcal{V},\mathcal{G})$, where $\tilde{\mathcal{G}}$ and $\mathcal{V}$ are separable Hilbert spaces. Assume that ${K_{0}}$ is of trace class, that Assumption D.5 holds, that $L_{n}=K(\cdot,x_{n})^{\star}$, $n\in{\mathbb{N}_{>0}}$, and that $\eta$ is $1$-subgaussian. Then, there exists an isometry $\iota_{\mathcal{H}_{K}}:\mathcal{H}_{{K_{0}}}\otimes\mathcal{V}\to\mathcal{H}_{K}$ such that, for all $Q\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ self-adjoint, positive semi-definite, and trace-class, and $\tilde{V}\in\mathcal{L}(\mathcal{H}_{{K_{0}}})$ self-adjoint, positive definite, diagonal, and with bounded inverse, it holds for all $\delta\in(0,1)$ that
@@
This follows immediately from Theorem D.14, since Proposition D.12 guarantees that its assumptions hold. ∎
@@
Corollary D.7 (Corollary D.20, RKHS case).
Under the setup and assumptions of Theorem D.6, do not assume that $\eta$ is $1$-subgaussian, but rather that there exists $\rho\in\mathbb{R}_{>0}$ and $R_{\mathcal{V}}\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ self-adjoint, positive semi-definite, and trace-class, such that $\eta$ is $R$-subgaussian, where $R:=\rho^{2}\iota_{\mathcal{G}}(\mathrm{id}_{\tilde{\mathcal{G}}}\otimes R_{\mathcal{V}})\iota_{\mathcal{G}}^{-1}$. Then, for all $\delta\in(0,1)$,
@@
Corollary D.8 (Corollary D.20, RKHS with trace-class kernel case).
Under the setup and assumptions of Theorem D.6, assume that $K$ itself is of trace class. Also, do not assume that $\eta$ is $1$-subgaussian, but rather that it is $\rho$-subgaussian for some $\rho\in\mathbb{R}_{>0}$. Then, we can take ${K_{0}}=K$, $\tilde{\mathcal{G}}=\mathcal{G}$, $\mathcal{V}=\mathbb{R}$, and $Q=\mathrm{id}_{\mathbb{R}}$ in Theorem D.6. In particular, for all $\delta\in(0,1)$,
@@
Proofs of Corollaries D.7 and D.8.
These results follow immediately from Corollary D.20. ∎
The next two sections are dedicated to stating and proving Theorem D.14 and Corollary D.20, which generalize Theorem D.6 and its corollaries to the case where $\mathcal{H}$ is not necessarily \@iaciRKHS RKHS. It relies on assuming the existence of the aforementioned appropriate factorization of $\mathcal{H}$ and of the operators $(L_{n})_{n}$, similarly as in the case of \@iaciUBD UBD kernel. We thus begin by formalizing the assumption.
D.3.2 Tensor factorization of Hilbert spaces and coefficients transformations
We have identified that, in the case $\mathcal{H}=\mathcal{H}_{k\cdot\mathrm{id}_{\mathcal{G}}}$, the core assumption for Theorem D.6 was that both $\mathcal{H}$ was the closure of the span of vectors of the form $f\cdot g$, with $f\in\mathcal{H}_{k}$ and $g\in\mathcal{G}$, and that $L_{n}$ could be identified with $\tilde{L}_{n}\otimes\mathrm{id}_{\mathcal{G}}$. This is the idea we generalize with the construction of this section. Specifically, we rigorously introduce tensor factorizations of Hilbert spaces, and formalize the assumption we make on the operators $(L_{n})_{n}$: they should have an appropriate form in the chosen factorization.
Definition D.9.

### v2

@@ Appendix C Summary of algorithms
Case 1 in Theorem 4.3 generalizes Corollary 3.4 in Abbasi-Yadkori [12] to outputs in separable Hilbert spaces without introducing conservatism, as it recovers exactly the same bound as in this reference under the assumptions there. It also generalizes Theorem 1 in Chowdhury and Gopalan [20] by removing the regularity assumptions connecting the input set $\mathcal{X}$ and the kernel $K$ (in the reference, $\mathcal{X}$ is compact and $K$ is bounded). It further generalizes this same result in yet another way by removing the assumption that $K$ should be trace-class, replacing it by the (significantly milder) assumption that $K_{0}$} should be so and hereby allowing the special case of a diagonal kernel $K=k\cdot\mathrm{id}_{\mathcal{G}}$. To the best of our knowledge, this is the first time-uniform uncertainty bound for KRR allowing this essential case.
@@
The assumption that the kernel $K$ is UBD in Case 1 above is the central methodological contribution of the proof. It enables us to generalize the proof method of Abbasi-Yadkori [12] by leveraging a tensor factorization of the RKHS $\mathcal{H}_{K}$ (Proposition E.12). Specifically, Proposition E.12 allows avoiding the usual problems encountered with non-trace-class kernels — more precisely, that trace-dependent quantities explode — by replacing them with the corresponding quantities for the elementary block $K_{0}$}. Intuitively, all of the “estimation machinery” is performed in the RKHS of $K_{0}$}, which is a trace-class kernel and for which classical methods work well, and we simply need to ensure that the noise is “compatible” with this structure. Here, this compatibility takes the form of the operator $R_{\mathcal{V}}$ being trace-class. Interestingly, a special case of Proposition E.12 has been used in Li et al. [41] (Theorem 1) to identify $\mathcal{H}_{K}$ with a space of Hilbert-Schmidt operators in the case $K=k\cdot\mathrm{id}_{\mathcal{G}}$, hereby avoiding problems with the trace of $K$ by working with that of $k$, similarly to us. While Li et al. [41] is focused on learning rates for conditional mean embeddings, the subsequent work [42] investigates (in some settings minimax optimal) learning rates for vector-valued KRR, utilizing essentially the same strategy of identifying KRR using a kernel of the form $K=k\cdot\mathrm{id}_{\mathcal{G}}$ with regularized estimation of Hilbert-Schmidt operators. Conceptually, both [41, 42] and our work use a tensor space construction to avoid problems with quantities related to trace-class operators, however, the technical details significantly differ. First, [41, 42] uses an identification with Hilbert-Schmidt operators, whereas we work with more general tensor product structures, that allow us formulate our results using UBD kernels. Second, [41, 42] aim at learning rates in a typical supervised learning setup, in particular, with random covariates, whereas we are interested in uniform (in the covariates) confidence bounds, and hence the proof techniques are rather different. Finally, the regularized least-squares algorithm analyzed by [41, 42] is not affected by whether the used kernel is trace-class or not, only the analysis needs to take this into account. In contrast, we need the tensor structure as formalized by our UBD assumption to even have a well-defined expression for the confidence bound due to appearance of the Fredholm determinant. We would like to remark that it appears that many arguments of Li et al. [41, 42] also hold under the more general assumption that $K$ is UBD. An interesting avenue for future work is thus to investigate to what extent this generalization carries over, and whether the derived learning rates, e.g., [41, Theorem 2], hold under UBD. Summarizing, we believe the UBD assumption and the resulting structure of the RKHS identified in Proposition E.12 is a relevant new idea to handle a large class of non-trace-class kernels and can be used in other areas of kernel methods such as learning theory with loss functions compatible with the structure.
@@
Appendix E Generalized confidence bounds in vector-valued least squares
We provide in this section the proof of Theorem 4.3. Specifically, we show a more general result which does not assume that the space $\mathcal{H}_{K}$ in which the estimator lies is \@iaciRKHS RKHS; only that it is a generic Hilbert space $\mathcal{H}$. The result we show is given later in this section in Theorem E.23 and pertains to confidence bounds in vector-valued least squares estimation under general assumptions on the sampling operators; KRR is then obtained for a specific choice of the measurement operators and Hilbert space where estimation is performed.
This section is organized as follows. We first provide in Section E.1 the general setup of vector-valued least squares estimation, and identify that the main challenge for confidence bounds is bounding a specific term involving the noise. We then provide in Sections E.2 and E.3 methods to bound this term under different assumptions that we formalize. The main intermediate results of those sections are Lemma E.4, Theorem E.14, and Corollary E.20, and are essential in the proof of Theorem E.23, which we only state and prove in Section E.4 as we provide a gentle introduction to its assumptions first. We finally leverage it in Section E.5 to show Theorem 4.3.
E.1 Setup and goal
@@
$\hskip 28.45274pt+\left\|\left(\sum_{n=1}^{N}L_{n}^{\ast}L_{n}+\lambda\mathrm{id}_{\mathcal{H}}\right)^{-1}\sum_{n=1}^{N}L_{n}^{\ast}y_{n}\right\|_{\sum_{n=1}^{N}L_{n}^{\ast}L_{n}+\mathrm{id}_{\mathcal{H}}}^{2}$
$\hskip 28.45274pt-\left\|\left(\sum_{n=1}^{N}L_{n}^{\ast}L_{n}+\lambda\mathrm{id}_{\mathcal{H}}\right)^{-1}\sum_{n=1}^{N}L_{n}^{\ast}y_{n}\right\|_{\sum_{n=1}^{N}L_{n}^{\ast}L_{n}+\lambda\mathrm{id}_{\mathcal{H}}}^{2}$
@@
[…]
@@
Our general goal is to obtain a finite-sample confidence region for the estimator $h_{N,\lambda}$ of $h^{\star}$. While we only state the precise result in Section E.4, we want — in short — a high-probability bound on the quantity $\lVert\bar{L}h_{N,\lambda}-\bar{L}h^{\star}\rVert_{\bar{\mathcal{G}}}$. Here $\bar{\mathcal{G}}$ is a Hilbert space and $\bar{L}\in\mathcal{L}_{\mathrm{b}}(\mathcal{H},\bar{\mathcal{G}})$ is a bounded operator capturing an appropriate property we are interested in; they are degrees of freedom of the problem. We begin with the following observation. This result, which is well-known in the case of KRR (see for example Abbasi-Yadkori [12, Chapter 3] or Chowdhury and Gopalan [19]), results from elementary algebraic manipulations, the triangle inequality, and properties of the operator norm.
Lemma E.2.
@@
$\begin{split}&\lVert\bar{L}(M_{N}^{\star}M_{N}+\lambda\mathrm{id}_{\mathcal{H}})^{-1/2}\rVert_{\mathcal{L}_{\mathrm{b}}(\mathcal{H},\bar{\mathcal{G}})}\\ &=\sqrt{\left\lVert\bar{L}(M_{N}^{\star}M_{N}+\lambda\mathrm{id}_{\mathcal{H}})^{-1/2}\left(\bar{L}(M_{N}^{\star}M_{N}+\lambda\mathrm{id}_{\mathcal{H}})^{-1/2}\right)^{\star}\right\rVert_{\mathcal{L}_{\mathrm{b}}(\mathcal{H})}}\\ &=\sqrt{\frac{1}{\lambda}\left\lVert\bar{L}\left(\mathrm{id}_{\mathcal{H}}-(M_{N}^{\star}M_{N}+\lambda\mathrm{id}_{\mathcal{H}})^{-1}M_{N}^{\star}M_{N}\right)\bar{L}^{\star}\right\rVert_{\mathcal{L}_{\mathrm{b}}(\mathcal{H})}}.\end{split}$ (21)
@@
E.2 Noise bound for deterministic sampling
@@
Assumption E.3 (Deterministic sampling).
@@
E.3 Noise bound for online sampling with coefficients-transforming sampling operators
@@
Assumption E.5 (Online sampling).
@@
We address this correlation by showing a self-normalized concentration inequality for
$\lVert S_{N}\rVert_{(M_{N}^{\star}M_{N}+\lambda\mathrm{id}_{\mathcal{H}})^{-1}}.$
The inequality is also time-uniform, meaning that the high-probability set on which the bound holds does not depend on the time $N\in{\mathbb{N}_{>0}}$. It is a generalization of Theorem 3.4 (and Corollary 3.6) in Abbasi-Yadkori [12] from $\mathcal{G}=\mathbb{R}$ to a general separable Hilbert space $\mathcal{G}$. These results are based on the method of mixtures [43]. While we mostly follow the proof strategy from Abbasi-Yadkori [12], we introduce a general structural assumption on the sampling operators (Definition E.11 and Proposition E.12) to handle the case where the sampling operators are not trace-class (in the RKHS case, the case where the kernel is not trace-class). Furthermore, in contrast to Abbasi-Yadkori [12] we do not employ an explicit stopping-time construction to obtain time uniformity, but rather use Ville’s inequality as in Whitehouse et al. [21].
E.3.1 Challenge and goal
@@
Our proof is based on the method of mixtures of Abbasi-Yadkori [12]. A close look at the proof of Theorem 3.4 in the reference reveals that the problematic determinant (24) arises when integrating the function $G_{N}$ (in the notations of the reference) against a Gaussian measure on $\mathcal{H}_{K}$. We circumvent this integration by leveraging a factorization of $\mathcal{H}_{K}$ as a tensor space. While such a factorization $\mathcal{H}_{K}\cong H_{1}\otimes H_{2}$ trivially exists for choices of separable $H_{1}$ and $H_{2}$ based on dimensional arguments (i.e., $\dim(\mathcal{H}_{K})=\dim(H_{1})\cdot\dim(H_{2})$), we choose one in which all of the sampling operators $L_{n}$, $n\in{\mathbb{N}_{>0}}$, admit themselves the simple form $L_{n}\cong\tilde{L}_{n}\otimes\mathrm{id}_{H_{2}}$, and for which $\tilde{L}_{n}$ has nice properties (specifically, $\tilde{L}_{n}^{\star}\tilde{L}_{n}$ is trace-class). In the case of a diagonal kernel — that is, $K=k\cdot\mathrm{id}_{\mathcal{G}}$ and $L_{n}=K(\cdot,x_{n})$, this construction is known and natural: the RKHS $\mathcal{H}_{K}$ is the (closure of the) span of vectors of the form $f\cdot g$, where $f\in\mathcal{H}_{k}$ and $g\in\mathcal{G}$; see for instance Paulsen and Raghupathi [35, Exercise 6.3]. Then, for any $h=\sum_{m=1}^{\infty}f_{m}\cdot g_{m}\in\mathcal{H}_{k\cdot\mathrm{id}_{\mathcal{G}}}$, introducing $h_{\otimes}=\sum_{m=1}^{\infty}f_{m}\otimes g_{m}\in\mathcal{H}_{{K_{0}}}\otimes\mathcal{G}$,
@@
In the more general case where $K$ is only UBD instead of diagonal, we prove that such a factorization also exists (Proposition E.12). In any case, this factorization allows us to write $\mathcal{H}_{K}\cong\mathcal{H}_{{K_{0}}}\otimes\mathcal{V}$. This enables evaluating (a modification of) the function $G_{N}$ on pure tensors $\tilde{h}\otimes v$, where $\tilde{h}\in\mathcal{H}_{{K_{0}}}$ and $v\in\mathcal{V}$. We then leverage a double mixture argument: one for the variable $\tilde{h}$, and one for the variable $v$. The one on $\tilde{h}$ is the same as in Abbasi-Yadkori [12]: a Gaussian measure on $\mathcal{H}_{{K_{0}}}$, which outputs the desired determinant involving $K_{0}$}. The one on $v$ is not Gaussian and enables recovering the norm we wish to bound, $\lVert S_{N}\rVert_{(V_{N}+\lambda\mathrm{id}_{\mathcal{H}_{K}})^{-1}}$ (up to some regularization term), when combined with Jensen’s inequality. Our double mixture strategy is fundamentally different from what is used for Chowdhury and Gopalan [19, Theorem 1], since we use two different types of mixture distributions, and the second mixture enters in a multiplicative manner (instead of an additive one). We can then conclude using standard arguments involving Ville’s inequality for positive supermartingales, similarly to Whitehouse et al. [21]. Summarizing, this method enables showing the following result, which is a specialization of the main outcome of this section to the case $\mathcal{H}=\mathcal{H}_{K}$ and $L_{n}=K(\cdot,x_{n})^{\star}$.
Theorem E.6 (Theorem E.14, RKHS case).
Under the notations of Sections E.1 and E.3.1, assume that $\mathcal{H}=\mathcal{H}_{K}$, where $K$ is UBD with block decomposition ${K_{0}}:\mathcal{X}\times\mathcal{X}\to\tilde{\mathcal{G}}$ and isometry $\iota_{\mathcal{G}}\in\mathcal{L}_{\mathrm{b}}(\tilde{\mathcal{G}}\otimes\mathcal{V},\mathcal{G})$, where $\tilde{\mathcal{G}}$ and $\mathcal{V}$ are separable Hilbert spaces. Assume that $K_{0}$} is of trace class, that Assumption E.5 holds, that $L_{n}=K(\cdot,x_{n})^{\star}$, $n\in{\mathbb{N}_{>0}}$, and that $\eta$ is $1$-subgaussian. Then, there exists an isometry $\iota_{\mathcal{H}_{K}}:\mathcal{H}_{{K_{0}}}\otimes\mathcal{V}\to\mathcal{H}_{K}$ such that, for all $Q\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ self-adjoint, positive semi-definite, and trace-class, and $\tilde{V}\in\mathcal{L}(\mathcal{H}_{{K_{0}}})$ self-adjoint, positive definite, diagonal, and with bounded inverse, it holds for all $\delta\in(0,1)$ that
@@
This follows immediately from Theorem E.14, since Proposition E.12 guarantees that its assumptions hold. ∎
@@
Corollary E.7 (Corollary E.20, RKHS case).
Under the setup and assumptions of Theorem E.6, do not assume that $\eta$ is $1$-subgaussian, but rather that there exists $\rho\in\mathbb{R}_{>0}$ and $R_{\mathcal{V}}\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ self-adjoint, positive semi-definite, and trace-class, such that $\eta$ is $R$-subgaussian, where $R:=\rho^{2}\iota_{\mathcal{G}}(\mathrm{id}_{\tilde{\mathcal{G}}}\otimes R_{\mathcal{V}})\iota_{\mathcal{G}}^{-1}$. Then, for all $\delta\in(0,1)$,
@@
Corollary E.8 (Corollary E.20, RKHS with trace-class kernel case).
Under the setup and assumptions of Theorem E.6, assume that $K$ itself is of trace class. Also, do not assume that $\eta$ is $1$-subgaussian, but rather that it is $\rho$-subgaussian for some $\rho\in\mathbb{R}_{>0}$. Then, we can take ${K_{0}}=K$, $\tilde{\mathcal{G}}=\mathcal{G}$, $\mathcal{V}=\mathbb{R}$, and $Q=\mathrm{id}_{\mathbb{R}}$ in Theorem E.6. In particular, for all $\delta\in(0,1)$,
@@
Proofs of Corollaries E.7 and E.8.
These results follow immediately from Corollary E.20. ∎
The next two sections are dedicated to stating and proving Theorem E.14 and Corollary E.20, which generalize Theorem E.6 and its corollaries to the case where $\mathcal{H}$ is not necessarily \@iaciRKHS RKHS. It relies on assuming the existence of the aforementioned appropriate factorization of $\mathcal{H}$ and of the operators $(L_{n})_{n}$, similarly as in the case of \@iaciUBD UBD kernel. We thus begin by formalizing the assumption.
E.3.2 Tensor factorization of Hilbert spaces and coefficients transformations
We have identified that, in the case $\mathcal{H}=\mathcal{H}_{k\cdot\mathrm{id}_{\mathcal{G}}}$, the core assumption for Theorem E.6 was that both $\mathcal{H}$ was the closure of the span of vectors of the form $f\cdot g$, with $f\in\mathcal{H}_{k}$ and $g\in\mathcal{G}$, and that $L_{n}$ could be identified with $\tilde{L}_{n}\otimes\mathrm{id}_{\mathcal{G}}$. This is the idea we generalize with the construction of this section. Specifically, we rigorously introduce tensor factorizations of Hilbert spaces, and formalize the assumption we make on the operators $(L_{n})_{n}$: they should have an appropriate form in the chosen factorization.
Definition E.9.